OpenStudy (anonymous):

How do i find the axis of symmetry of y–2=2(x +1)^2?

OpenStudy (saifoo.khan):

y–2=2(x +1)^2? y = 2(x+1) + 2 solve the inner parenthesis for x. (x+1) = 0 x = 1 <--- axis of symmetry.

OpenStudy (anonymous):

where did the ^2 go?

OpenStudy (saifoo.khan):

that remains there. it is not needed here.

OpenStudy (anonymous):

okay so woulndt it be x= -1?